1. A woman has 4 children. The eldest is 3 years older than the second child, who is 3 years older than the youngest child. The youngest child is one half the age of the oldest child. How old are the children.
2. A card is cut from a pack of playing cards. Determine the probability that it will be an ace or the king of hearts.
3. A coin is tossed and a die is rolled. What is the probability of getting:
· A tail and a 5?
· A head and an even number?

Answer by stanbon(48568)

1. A woman has 4 children. The eldest is 3 years older than the second child, who is 3 years older than the youngest child. The youngest child is one half the age of the oldest child. How old are the children.
---------
Let the youngest child be age "x".
Then oldest is age "2x".
The second child is "2x-3"
EQUATION:
2nd is 3 years older than youngest
2x-3=x+3
x=6 (age of youngest)
2x=12 (oldest)
2x-3=9 (second oldest child)
=====================
2. A card is cut from a pack of playing cards. Determine the probability that it will be an ace or the king of hearts.
Prob(ace or king of hearts)= Prob(ace)+Prob(king of hears)=4/52+1/52=5/52
================
3. A coin is tossed and a die is rolled. What is the probability of getting:
· A tail and a 5?
Prob(tail) * Prob(five) = (1/2) * (1/6) = 1/12
· A head and an even number?
Prob(head) * Prob(even number) = (1/2) * (3/6) = 1/4
==========
Cheers,
Stan H.